OpenStudy (anonymous):

express 2x + 2y = 6 in point-slope form.

OpenStudy (anonymous):

Suppose x=1, so at that position the y value is 2. This gives us a point on the line of (1,2). By rearranging the equation, we get y=-x+3, then m=-1. Use the point/slope form to get y-2=-1(x-1).

OpenStudy (anonymous):

Note that you could use any other point on the line, as well.

OpenStudy (anonymous):

why is it -(x-1)?

OpenStudy (anonymous):

The formula is y-y1=m(x-x1). In this case, we calculated m=-1. In simpler notation, it can be written -(x-1).